OpenStudy (anonymous):

Hiram raises earthworms. In a square of compost 4 ft by 4 ft, he can have 1,000 earthworms. How many earthworms can he have if his square of compost has a side length that is 7 times longer? How do i solve this!?

OpenStudy (cwrw238):

the ratios of the areas of small to larger is the ratio of the squares of the sides so as new side is 4 * 5 = 20 ft ratio of their areas is 4^2 : 20^2 = 16 : 400 or 1 : 25

OpenStudy (cwrw238):

so the area is 25 times larger and now you just multiply ...

OpenStudy (anonymous):

THANX:) but multiply what though?

OpenStudy (cwrw238):

1000 by 25

OpenStudy (anonymous):

but that is not one of the answer choices...the answerchoices are a7000 b28,000 c196,000 d49,000

OpenStudy (cwrw238):

lets do it another way area of small square = 4*4 = 16 sq ft the side of big square is 5 times longer that smaller so its 4*5 = 20 ft and its area is 20*20 = 400 sq ft so its 400 / 16 = 25 times bigger number of earthworms = 25 * 1000 = 25,000

OpenStudy (cwrw238):

either the answers are all wrong or maybe you've got the question wrong

OpenStudy (anonymous):

IM SOO SORRRY i had the question wrong lol its 7 instead of 5 Sorry

OpenStudy (cwrw238):

small area = 4*4 = 16 large area = (4*7)^2 = 28^2 so ratio is 28* 28 / 16 use your calculator to find this and then multiply by 1000

OpenStudy (anonymous):

so 1000 x 28 ?